OpenStudy (anonymous):

Two end points of a line segment are (-12, -2) and (-6, -10). What are the coordinates of the point on the line through which its bisector passes? (-18, -12) (-9, -6) (-7, -8) (-11, -4)

ganeshie8 (ganeshie8):

hint : bisector goes thru midpoint of the segment.

OpenStudy (anonymous):

Still Lost...

ganeshie8 (ganeshie8):

use the midpoint formula

ganeshie8 (ganeshie8):

midpoint of (-12, -2) and (-6, -10) : \(\large (\frac{-12-6}{2}, \frac{-2-10}{2} )\) = ?

OpenStudy (anonymous):

Thanks!!

ganeshie8 (ganeshie8):

np :)